OpenStudy (anonymous):

Calculate the period (T) of uniform circular motion if the velocity is 40.0 m/s and centripetal acceleration is 20.0 m/s2.

OpenStudy (anonymous):

@YanaSidlinskiy

OpenStudy (yanasidlinskiy):

Centripetal acceleration, a= v^2/r This gives r = v^2/a = 40.0^2/20.0 = 80.0m Angular velocity, w = 2pi/T. Also w = v/r. So 2pi/T = v/R 2pi/T = 40.0/80.0 = 0.5 T = 2pi/0.5 = 12.6s

OpenStudy (anonymous):

A bit confusing but i understand it a little more thank you

OpenStudy (yanasidlinskiy):

What's confusing?

OpenStudy (anonymous):

Just the process of getting to it

OpenStudy (yanasidlinskiy):

Notice, That I have seperated them. One is Centripetal acceleration and the other one is Angular velocity. To get to the answer, we would add.

OpenStudy (anonymous):

Okay okay it makes more sense now

OpenStudy (yanasidlinskiy):

:D

OpenStudy (anonymous):

Thanks for helping your awesome :D

OpenStudy (yanasidlinskiy):

Your Welcome:D I wouldn't say I am though..